### v2

@@ Abstract
We study particle systems interacting via hitting times on sparsely connected graphs, following the framework of Lacker, Ramanan and Wu [28]. We provide general robustness conditions that guarantee the well-posedness of physical solutions to the dynamics, and demonstrate their connections to the dynamic percolation theory. We then analyze the limiting behavior of the particle systems, establishing the continuous dependence of the joint law of the physical solution on the underlying graph structure with respect to local convergence and studying the convergence of the global empirical measure, which extends the general results by Lacker et al. to systems with singular interactions. The model proposed provides a general mathematical framework in continuous time for analyzing systemic risks in large sparsely connected financial networks with a focus on local interactions, featuring instantaneous default cascades.
@@ 1 Introduction
We study the following interacting particle system on a countable and locally finite graph $G=(V,E)$:
$X_{v}(t)$ $=x_{v}+B_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t}\}},\quad v\in V$ (1.1)
$D_{t}$ $=\{v\in V\mathrel{\mathop{\ordinarycolon}}\,\inf_{s\in[0,t]}X_{v}(s)\leq 0\},$ (1.2)
which can be seen as a stylized model for an inter-connected network of banks with mutual lending, in which the default of a bank leads to immediate losses to its creditors. In this system, each vertex $v\in V$ represents a particle (financial institution, such as a bank) with initial healthiness (asset value) $x_{v}\in[0,\infty)$, $(B_{v})_{v\in V}$ are a collection of independent $V$-indexed Brownian motions driving the dynamics, $N_{G}^{-}(v)\mathrel{\mathop{\ordinarycolon}}=\{u\in V\mathrel{\mathop{\ordinarycolon}}(u,v)\in E\}$ denotes the in-neighborhood of $v$, and $c_{uv}\geq 0$ is the loss suffered by particle $v$ if particle $u$ dies (defaults), i.e. the healthiness of particle $u$ gets as low as 0. The set-valued process $t\mapsto D_{t}$ records the set of banks which default no later than time $t$.
@@
In this model, when a bank defaults, it triggers immediate losses to its neighbors, after which it cannot have a second impact on the system. In that sense, the defaulted bank can be treated as “removed” from the network, even though the network structure remains the same and the corresponding driving noise $B_{v}(t)$ continues to evolve in a purely imaginary manner.
The above system can be formulated as a fixed-point problem. Given a realization of the graph $G$ and the driving noises $(B_{v})_{v\in V}$, define the operator $\Gamma$, which maps a set-valued process $D=(D_{t})_{t\geq 0}$ to another set-valued process $\Gamma[D]=(\Gamma[D]_{t})_{t\geq 0}$ via
$\Gamma[D]_{t}\mathrel{\mathop{\ordinarycolon}}=\Big{\{}v\in V\mathrel{\mathop{\ordinarycolon}}\,\inf_{s\in[0,t]}\Big{(}x_{v}+B_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t}\}}\Big{)}\leq 0\Big{\}}.$
@@
We first point out that equations (1.1) and (1.2) do not uniquely pin down a dynamics, even in the simplest case:
@@
Let $G=(V=\{0,1\},E=\{(0,1),(1,0)\})$, $x_{0}=x_{1}=1$ and $c_{01}=c_{10}=1$. Then,
  1. 1.
    there exists a solution $(X,D)$ such that $D_{t}=\emptyset$ for all $t<\tau$, where $\tau\mathrel{\mathop{\ordinarycolon}}=\inf\{t\geq 0\mathrel{\mathop{\ordinarycolon}}\,B_{0}(t)\leq-1\,\,\text{or}\,\,B_{1}(t)\leq-1\}$.
  2. 2.
@@
  • (a)
    The map $t\mapsto D_{t}$ is right-continuous,
  • (b)
@@
    $D_{t}^{(0)}$ $\mathrel{\mathop{\ordinarycolon}}=\{v\in V\mathrel{\mathop{\ordinarycolon}}\inf_{s\in[0,t)}X_{v}(s)\leq 0\},$
    $D_{t}^{(N+1)}$ $\mathrel{\mathop{\ordinarycolon}}=D_{t}^{(N)}\cup\Big{\{}v\in V\mathrel{\mathop{\ordinarycolon}}\,x_{v}+B_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t}^{(N)}\}}\leq 0\Big{\}},$
@@
  • (3)
    As we do not rule out potential defaults at $t=0$, we need to extend the time range for the dynamics to $1,\infty)$ by setting
    $X_{v}(t)=x_{v},\quad D_{t}=\emptyset,\quad B_{v}(t)=0,\quad t\in[-1,0).$
    With this extension, it makes sense to talk about the left limits of the dynamics at $t=0$, which encode the initial conditions for the dynamics. Physical solution $(X,D)$, if it exists, will have all of its paths $X_{v}(\cdot)$ and $\mathbf{1}_{\{v\in D_{\cdot}\}}$ belong to the space $\mathcal{D}\mathrel{\mathop{\ordinarycolon}}=D([-1,\infty))$. A brief introduction to the space $\mathcal{D}$ as well as the $M_{1}$ topology on it is given in Section 2.4.
@@
It follows immediately from the definition that the minimal solution, if it exists, must be unique. The technique of identifying physical solutions based on minimal solutions originates from [13].
The well-posedness of the system poses several challenges when the underlying graph $G$ is infinite. First, if the Brownian motions $(B_{v})_{v\in V}$ are i.i.d., the zero-hitting times are typically dense in the time axis $(0,\infty)$. As a result, the system may lack a strict separation between continuous and jump regimes. Second, without further assumptions, it may not be guaranteed that the system can remain stable, even for a sufficiently short time, as illustrated by the following example.
@@
We take $V\mathrel{\mathop{\ordinarycolon}}=\mathbb{N}=\{0,1,2,$...\} and $E\mathrel{\mathop{\ordinarycolon}}=\{(n,n+1)\mathrel{\mathop{\ordinarycolon}}\,n\in\mathbb{N}\}$ with $c_{n+1,n}=1$. Let $x_{n}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{2^{n+1}}$ for all $n\geq 0$ and let $(B_{n})_{n\geq 0}$ be independent standard Brownian motions. Then any solution $(X,D)$ to equations (1.1) and (1.2) satisfies $D_{0+}=V=\mathbb{N}$ with probability $1$. In particular, there is no physical solution to equations (1.1) and (1.2) with the initial conditions given above.
To ensure the well-posedness of the system, we introduce a set of robustness assumptions on the problem configuration.
@@
The initial configuration $(G,c,x)$ is said to induce $\delta$-robust systems if, with probability one, it holds that for every $t\geq 0$, the set
$\Big{\{}v\in V\mathrel{\mathop{\ordinarycolon}}\exists s\in[t,t+\delta],\text{ s.t. }0\leq x_{v}+B_{v}(s)\leq\sum_{u\in N_{G}^{-}(v)}c_{uv}\Big{\}}\,$
contains no infinite weakly connected component, where $(B_{v})_{v\in V}$ are independent standard Brownian motions conditioned on $(G,c,x)$. The initial configuration $(G,c,x)$ is said to induce robust systems if it drives $\delta$-robust systems with $\delta=0$.
A collection of sufficient conditions for $\delta$-robustness are provided in Lemma 3.5. As we will show, $\delta$-robustness is sufficient for the well-posedness of physical solutions to equations (1.1) and (1.2).
@@
For each initial configuration $(G,c,x)\in\mathcal{G}_{*}[\mathbb{R}]$ that induces $\delta$-robust systems, there exists a unique physical solution to the equations (1.1) and (1.2). Moreover, this solution coincides with the minimal solution.
@@
    $\mu^{n}=\frac{1}{|G_{n}|}\sum_{v\in G_{n}}\delta_{(X_{v}^{n},\tau_{v}^{n})}$
@@
If $(G_{n},c^{n},x^{n})$ is a sequence of initial configurations that induce robust systems and that $\mathcal{L}(G_{n},c^{n},x^{n})\to\mathcal{L}(G,c,x)$ in $\mathcal{P}(\mathcal{G}_{*}[\mathbb{R}])$ as $n\to\infty$ where the limit configuration $(G,c,x)$ induces a $\delta$-robust system, then $\mathcal{L}(G_{n},X^{n},D^{n})\to\mathcal{L}(G,X,D)\,\text{in }\mathcal{P}(\mathcal{G}_{*}[\mathcal{D}^{2}]),$ where $(X^{n},D^{n})$ and $(X,D)$ denote the unique physical solutions associated with $(G_{n},c^{n},x^{n})$ and $(G,c,x)$, respectively.
@@
If $(G_{n},c^{n},x^{n})$ is a sequence of initial configurations such that each $G_{n}$ is finite, and suppose that $(G_{n},c^{n},x^{n})$ converges in probability in the local weak sense to $(G,c,x)$, which is a random element in $\mathcal{G}_{*}[\mathbb{R}]$ that induces a $\delta$-robust system. Then $(G_{n},X^{n},D^{n})$ converges in probability in the local weak sense to $(G,X,D)$. In particular, the empirical measure $\mu^{n}$ converges in probability to $\mathcal{L}(X_{o},\tau_{o})$, where $o$ is the root of $G$.
Remark 1.11.
Although we assumed that $(B_{v})_{v\in V}$ are independent Brownian motions when introducing the model and the main results, our theory actually applies to a much more general class of noise processes, as illustrated by Lemma 3.5, Theorem 3.19 and Theorem 3.21, in which we essentially treat the noise processes $(B_{v})_{v\in V}$ as part of the inputs to the dynamics. This extent of abstractness not only brings convenience for the mathematical analysis of the system, but also supports more realistic models for future investigations.
The main technical challenge in proving Theorems 1.8 and 1.9 lies in establishing qualitative control of the “propagation speed” of default cascades. Although the interactions in (1.1)–(1.2) are local, it remains possible in principle for a cascade to spread instantaneously across a large portion of the network. This phenomenon does not arise in [28] and cannot be handled by the techniques developed there. To overcome this difficulty, we analyze in detail the causality structure of the defaults and rigorously re-establish the locality of the interactions (Section 3.3), relying on tools from dynamic percolation theory. This step quantifies how the network topology constrains the growth of the default sets and forms the core technical contribution of the paper. By contrast, the proof of Theorem 1.10 is more direct than that of [28, Theorem 3.7], since the noise processes $(B_{v})_{v\in V}$ may be incorporated into the inputs of the dynamics.
The remainder of this section provides a brief review of related literature. In Section 2, we introduce the main definitions and notations used throughout the paper. Section 3 presents the main results and their proofs. The proof of Theorem 1.8 is established in Section 3.1 and 3.2. Theorem 1.9 is proved in Section 3.4, and Theorem 1.10 in Section 3.5. In Section 3.6, we demonstrate the connection between the model studied in this paper and other models of systemic risks in which losses are not realized immediately. Additional technical proofs are provided in the Appendix.
@@
To manage the complexity of analyzing continuous dynamics on large-scale financial networks, recent work, including [9, 10], has adopted mean-field formulations which aim to describe the macroscopic behavior of the system as the number of nodes tends to infinity by focusing on a representative particle–a phenomenon referred to as propagation of chaos, under assumptions of complete symmetry. This approach has proven fruitful in both stochastic control and game-theoretic settings. The pioneering works [15, 16] introduced the instantaneous contagion and the notion of physical solutions to mean-field particle systems and established the corresponding propagation of chaos. Also revealed are their deep connections, as a probabilistic method, to the supercooled Stefan problem, a singular free boundary problem, the well-posedness of which has remained challenging for decades. Following this framework, [17] established the global-in-time well-posedness of physical solutions and studied their regularities, and [13] proposed the notion of minimal solutions, pointed out their connections to physical solutions, and established the propagation of minimality. These models and their variants have been further developed and applied to the study of financial networks and neuronal networks as in [25, 31, 32, 24, 4, 26, 23, 30, 3, 14, 6]. It is also worth mentioning that there has been recent efforts as in [21, 22] towards analyzing how the spatial structure plays a role in cascades when the underlying space is continuous.
While mean field models have proven instrumental in describing the asymptotic behavior of large populations, they are often based on assumptions of full homogeneity and dense connectivity. The articles [32, 19] incorporated heterogeneity into the mean-field framework by constructing and analyzing multi-type particle systems, with the interactions between the types given by probability kernels. Within the types, the particles are still exchangeably distributed, densely connected, and each individual particle interacts with the whole distribution and has negligible effect on the rest of the system. Such assumptions can be too restrictive in the study of financial networks, which are typically sparsely connected and exhibit local heterogeneous interactions driven by direct bilateral exposures, as illustrated by empirical evidence given in [7, 11, 12]. To address this gap, we adopt the general theory developed in [28] of the local convergence for networks of interacting stochastic processes, which is particularly well-suited for modeling cascades in networks, where contagion spreads through direct neighboring exposure rather than global interaction.

### v3

@@ Abstract
We introduce a family of particle systems on sparse graphs where local interactions occur via hitting times, providing a dynamic and tractable model for default cascades in large sparsely-connected financial networks. Building on the framework of [26], we extend convergence theory to systems with singular interactions, capturing the abrupt and discontinuous nature of systemic events. We establish conditions for well-posedness through a minimality principle and connect fragility to dynamic percolation thresholds. Our analysis demonstrates continuity of the joint law of defaults with respect to local graph convergence, establishes convergence of empirical distributions, and characterizes the default time distribution in tree-like networks. This framework offers a rigorous and flexible foundation for modeling systemic risk in evolving financial systems, featuring continuous-time dynamics, heterogeneous and local interactions, and instantaneous default cascades.
@@ 1 Introduction
The global financial crisis of 2008 and the more recent bank failures in 2023 highlighted how interconnections among financial institutions can amplify idiosyncratic shocks into systemic events. This has made the analysis of default cascades in financial networks a central problem in systemic risk. Many existing models describe interbank contagion through static clearing mechanisms or discrete-time dynamics. Although these approaches have yielded important insights, they are less well suited to capturing the continuous-time, discontinuous, and self-exciting nature of contagion generated by sequential defaults. In this paper, we introduce a continuous-time locally interacting particle system for default cascades on networks, providing a tractable framework for the analysis of systemic fragility and resilience.
We propose to study the following interacting particle system on a countable and locally finite graph $G=(V,E)$:
$X_{v}(t)$ $=x_{v}+Z_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t}\}},\quad v\in V$ (1.1)
$D_{t}$ $=\big\{v\in V\mathrel{\mathop{\ordinarycolon}}\,\inf_{s\in[0,t]}X_{v}(s)\leq 0\big\}.$
It can be seen as a stylized model for an inter-connected network of banks with mutual lending, in which the default of a bank leads to immediate losses to its creditors. In this system, each vertex $v\in V$ represents a particle (financial institution, such as a bank) with initial health (asset value) $x_{v}\in[0,\infty)$, $(Z_{v})_{v\in V}$ are a pre-determined collection of $V$-indexed càdlàg stochastic processes (e.g. Brownian motions or more general Lévy processes) which represent the noise driving the dynamics, $N_{G}^{-}(v)\mathrel{\mathop{\ordinarycolon}}=\{u\in V\mathrel{\mathop{\ordinarycolon}}(u,v)\in E\}$ denotes the in-neighborhood of $v$, and $c_{uv}\geq 0$ is the loss suffered by particle $v$ if particle $u$ dies (or defaults) [^4] Rigorously speaking, the loss triggered by the default of a bank depends on multiple factors such as interbank lending relationships, asset and liability structure, counterparty risk in derivatives, clearing and settlement mechanism, investor sentiment, and regulatory scrutiny. We choose to set the default loss as a pre-determined constant to abstract away from the idiosyncratic factors that determine loss severity and concentrate on the structural dynamics of systemic risk transmission., that is, particle $u$’s health falls to 0. The set-valued process $t\mapsto D_{t}$ records the set of banks which default no later than time $t$.
@@
In this model, when a bank defaults, it triggers immediate losses to its neighbors, after which it cannot have a second impact on the system. In that sense, the defaulted bank can be treated as “removed” from the network, even though the network structure remains the same and the corresponding driving noise $Z_{v}(t)$ continues to evolve in a purely imaginary manner.
The above system can be formulated as a pathwise fixed-point problem. Given a realization of the graph $G$ and the driving noises $(Z_{v})_{v\in V}$, we define the operator $\Gamma$, which maps a set-valued process $D=(D_{t})_{t\geq 0}$ to another set-valued process $\Gamma[D]=(\Gamma[D]_{t})_{t\geq 0}$ via
$\Gamma[D]_{t}\mathrel{\mathop{\ordinarycolon}}=\Big\{v\in V\mathrel{\mathop{\ordinarycolon}}\,\inf_{s\in[0,t]}\Big(x_{v}+Z_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t}\}}\Big)\leq 0\Big\}.$
@@
We first point out that equations (1.1) do not uniquely pin down a dynamic, even in the simplest case:
@@
Let $G\mathrel{\mathop{\ordinarycolon}}=\Big(V=\{1,2\},E=\{(1,2),(2,1)\}\Big)$, $x_{1}=x_{2}=1$ and $c_{12}=c_{21}=1$. Then,
  1. (a)
    there exists a solution $(X,D)$ such that $D_{t}=\emptyset$ for all $t<\tau$, where $\tau\mathrel{\mathop{\ordinarycolon}}=\inf\{t\geq 0\mathrel{\mathop{\ordinarycolon}}\,Z_{1}(t)\leq-1\,\,\text{or}\,\,Z_{2}(t)\leq-1\}$.
  2. (b)
@@
  1. (a)
    the map $t\mapsto D_{t}$ is right-continuous, and
  2. (b)
@@
    $D_{t}^{(0)}$ $\mathrel{\mathop{\ordinarycolon}}=D_{t-}\cup\Big\{v\in V\mathrel{\mathop{\ordinarycolon}}\,x_{v}+Z_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t-}\}}\leq 0\Big\},$
    $D_{t}^{(N+1)}$ $\mathrel{\mathop{\ordinarycolon}}=D_{t}^{(N)}\cup\Big\{v\in V\mathrel{\mathop{\ordinarycolon}}\,x_{v}+Z_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t}^{(N)}\}}\leq 0\Big\},\quad N\geq 0,$
@@
  • (c)
    At the initial time $t=0$, we do not rule out potential discontinuities of the system (such as the occurrence of defaults), hence we need to extend the time range for the dynamics to $1,\infty)$ by setting
    $X_{v}(t)=x_{v},\quad D_{t}=\emptyset,\quad Z_{v}(t)=0,\quad t\in[-1,0).$
    With this extension, it becomes meaningful to consider the left limits of the dynamics at $t=0$, which encode the initial conditions for the dynamics. Physical solution $(X,D)$, if it exists, will have all of its paths $X_{v}(\cdot)$ and $\mathbf{1}_{\{v\in D_{\cdot}\}}$ belong to the space $\mathcal{D}\mathrel{\mathop{\ordinarycolon}}=D([-1,\infty))$. A brief introduction to the space $\mathcal{D}$ as well as the $M_{1}$ topology on it is given in Appendix A.
@@
It follows immediately from the definition that the minimal solution, if it exists, must be unique.The idea of characterizing physical solutions through minimal solutions goes back to [14].
Establishing the well-posedness of the system becomes challenging when the underlying graph $G$ is infinite. First, if the noise processes $(Z_{v})_{v\in V}$ are i.i.d., the zero-hitting times are typically dense in the time axis $(0,\infty)$. As a result, the system may lack a strict separation between continuous evolution and jump regimes. Second, without further assumptions, stability cannot be guaranteed even on arbitrarily short time intervals, as the following example demonstrates.
@@
We take $V\mathrel{\mathop{\ordinarycolon}}=\mathbb{N}=\{0,1,2,$...\} and $E\mathrel{\mathop{\ordinarycolon}}=\{(n,n+1)\mathrel{\mathop{\ordinarycolon}}\,n\in\mathbb{N}\}$ with $c_{n+1,n}=1$. Let $x_{n}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{2^{n+1}}$ for all $n\geq 0$ and let $(Z_{n}=B_{n})_{n\geq 0}$ be independent standard Brownian motions. Then any solution $(X,D)$ to equations (1.1) satisfies $D_{0+}=V=\mathbb{N}$ with probability 11. In particular, there is no physical solution to equations (1.1) and with the initial conditions given above.
To ensure the well-posedness of the system, we introduce a set of robustness conditions on the problem configuration.
@@
The problem configuration $(G,c,x,Z)$ is said to drive $\delta$-robust systems if, with probability one, it holds that for every $t\geq 0$, the set
$\Big\{v\in V\mathrel{\mathop{\ordinarycolon}}\exists s\in[t,t+\delta],\text{ s.t. }x_{v}+Z_{v}(s-)\geq 0\quad\text{and}\quad x_{v}+Z_{v}(s)\leq\sum_{u\in N_{G}^{-}(v)}c_{uv}\Big\}$
contains no infinite weakly connected component. The problem configuration $(G,c,x,Z)$ is said to drive robust systems if it drives $\delta$-robust systems with $\delta=0$.
A collection of sufficient conditions for $\delta$-robustness are provided in Lemma 3.5. As it turns out, $\delta$-robustness is sufficient for the well-posedness of physical solutions to equations (1.1). Before stating our main results, we state the general assumptions imposed on the driving noise component $Z$ throughout the paper:
Assumption 1.8.
  1. (a)
    $Z$ has no upward jumps:
    $Z(t)-Z(t-)\leq 0,\quad\forall t\geq 0.$
  2. (b)
    $Z$ satisfies the downward crossing property: for any fixed stopping time $\tau$,
    $\mathbb{P}\Big[\forall h>0,\,\inf_{s\in[0,h]}(Z(\tau+s)-Z(\tau))<0\Big]=1.$
Remark 1.9.
The functional that takes a càdlàg process to its zero-hitting time is continuous at those processes satisfying the downward crossing property (see e.g. [15, Proposition 5.8]). Formally speaking, any càdlàg stochastic process with a Brownian component satisfies the above downward crossing property. The scope of financial models encompassed by our framework is discussed further in Lemma 3.5, Remark 3.6, and Section 4.1.
@@
For each problem configuration $(G,c,x,Z)\in\mathcal{G}_{*}[\mathbb{R}\times\mathcal{D}]$ that drives $\delta$-robust systems, there exists a unique physical solution to the equations (1.1). Moreover, this solution coincides with the minimal solution.
@@
    $\mu^{n}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{|G_{n}|}\sum_{v\in G_{n}}\delta_{(X_{v}^{n},\tau_{v}^{n})}$
@@
Let $(G_{n},c^{n},x^{n},Z^{n})$ be a sequence of problem configurations that drive robust systems and that $\mathcal{L}(G_{n},c^{n},x^{n},Z^{n})\to\mathcal{L}(G,c,x,Z)$ in $\mathcal{P}(\mathcal{G}_{*}[\mathbb{R}\times\mathcal{D}])$ as $n\to\infty$, where the limit configuration $(G,c,x,Z)$ drives $\delta$-robust systems. Then $\mathcal{L}(G_{n},X^{n},D^{n})\to\mathcal{L}(G,X,D)\,\text{in }\mathcal{P}(\mathcal{G}_{*}[\mathcal{D}^{2}]),$ where $(X^{n},D^{n})$ and $(X,D)$ denote the unique physical solutions associated with $(G_{n},c^{n},x^{n},Z^{n})$ and $(G,c,x,Z)$, respectively.
@@
Let $(G_{n},c^{n},x^{n},Z^{n})$ be a sequence of problem configurations such that each each $G_{n}$ is finite, and suppose that $(G_{n},c^{n},x^{n},Z^{n})$ converges in probability in the local weak sense to a random element $(G,c,x,Z)\in\mathcal{G}_{*}[\mathbb{R}\times\mathcal{D}]$, which drives $\delta$-robust systems. Then the corresponding solutions $(G_{n},X^{n},D^{n})$ converges in probability in the local weak sense to $(G,X,D)$. In particular, the empirical measure $\mu^{n}$ converges in probability to $\mathcal{L}(X_{o},\tau_{o})$, where $o$ is the root of $G$.
The main technical challenge in proving Theorems 1.10 and 1.11 is to qualitatively control the “propagation speed” of default cascades. Although the interactions in the system (1.1) are local, it remains possible in principle for a cascade to spread instantaneously across a large portion of the network, exacerbating the discontinuity of the system under consideration. This phenomenon does not arise in [26] and therefore falls outside the scope of the methods developed there. To overcome this difficulty, our approach is to analyze carefully the causal structure of defaults and thereby re-establish locality of the dynamics, which is carried out in Section 3.3 using tools from dynamic percolation theory. This step quantifies how the network topology constrains the growth of the default sets and forms the core technical contribution of the paper. In contrast, the proof of Theorem 1.12 is more direct than that of [26, Theorem 3.7], since the noise processes $(Z_{v})_{v\in V}$ may be incorporated into the inputs of the dynamics.
The remainder of this section provides a brief review of related literature. In Section 2, we introduce the main definitions and notations used throughout the paper. Section 3 presents the main results and their proofs. The proof of Theorem 1.10 is established in Section 3.1 and 3.2. Theorem 1.11 is proved in Section 3.4, and Theorem 1.12 in Section 3.5. In Section 3.6, we demonstrate the connection between the model studied in this paper and other models of systemic risks in which losses are not realized immediately. In Section 4.1, we discuss potential generalization of our framework to include both major banks and local banks. Additional technical proofs are provided in the Appendices.
@@
To manage the complexity of analyzing continuous dynamics on large-scale financial networks, recent work, including [9, 10], has adopted mean-field formulations which aim to describe the macroscopic behavior of the system as the number of nodes tends to infinity by focusing on a representative particle, which is a phenomenon referred to as propagation of chaos, under assumptions of complete symmetry and weak interactions. This approach has proven fruitful in both stochastic control and game-theoretic settings. The pioneering works [17, 15] introduced the instantaneous contagion and the notion of physical solutions to mean-field particle systems and established the corresponding propagation of chaos. Also revealed are their deep connections, as a probabilistic method, to the supercooled Stefan problem, a singular free boundary problem, the well-posedness of which has remained challenging for decades. Following this framework, [16] established the global-in-time well-posedness of physical solutions and studied their regularities, and [14] proposed the notion of minimal solutions, pointed out their connections to physical solutions, and established the propagation of minimality. These models and their variants have been further developed and applied to the study of financial networks and neuronal networks as in [24, 29, 30, 22, 3, 23, 21, 28, 4, 13, 6].
While mean-field models have proven instrumental in describing the asymptotic behavior of large populations, they are often based on assumptions of full homogeneity and dense connectivity. The articles [30, 19] incorporated heterogeneity into the mean-field framework by constructing and analyzing multi-type particle systems, with the interactions between the types given by probability kernels. Within the types, the particles are still exchangeably distributed, densely connected, and each individual particle interacts with the whole distribution and has negligible effect on the rest of the system. Such assumptions can be too restrictive in the study of financial networks, which are typically sparsely connected and exhibit local heterogeneous interactions driven by direct bilateral exposures, as illustrated by empirical evidence given in [7, 11, 12]. To address this gap, we adopt the general theory developed in [26] of the local convergence for networks of interacting stochastic processes, which is particularly well-suited for modeling cascades in networks, where contagion spreads through direct neighboring exposure rather than global interaction. It is worth emphasizing that this form of large sparse graph limits is not covered by the theory of mean-field or graphon limits, as the correlation between the state processes of any fixed pair of vertices will typically not vanish in the limit.